please help me solve these equations using elimination y=(1/3)x+2 and y=(5/2)x-1

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Multiply either equation (BUT NOT BOTH) by -1. Add the two equations. Solve the result for . Using the value for , substitute into either original equation and then calculate the value of

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

solve these equations using elimination
y=(1/3)x+2
y=(5/2)x-1
-----
Subtract and solve for "x":
[(5/2)-(1/3)]x -3 = 0
----
[(15-2)/6]x = 3
-----
(13/6)x = 3
x = 18/13
-------
Solve for "y":
y = (1/3)(18/13)+2
-----
y = (6/13)+(26/13)
y = 32/13
------------------------
Cheers,
Stan H.